OpenStudy (anonymous):

will give medal graphing f(x)= 1/4x

OpenStudy (anonymous):

i know that its a reciprocal function i just dont understand what the four means

OpenStudy (anonymous):

basically i just need an explanation of the four. like does it compress or stretch, move or whatever

OpenStudy (anonymous):

4x is the denomiantor, or just 4?

OpenStudy (anonymous):

\[F(x)=\frac{1}{4x}\]this?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you want to graph it by hand?

OpenStudy (anonymous):

well i have like four choices and i just have to pick which one is correct.

OpenStudy (anonymous):

i just dont understand what the four does when im graphing

OpenStudy (anonymous):

well, this is what I can help you with: as x approaches infinity (saying to the very right) the y is going to be almost a zero. Because 1/(4 times a big number) = almost zero right?

OpenStudy (anonymous):



OpenStudy (anonymous):

it is going to approach zero very closely, but never actually hit zero.

OpenStudy (anonymous):

okay so would it b graphed the same as f(x)= 1/x orrrrrr

OpenStudy (anonymous):

same as f(x)=1/x BUT... when you multiply the x times 1/4 (the scale factor of a number, the absolute value of which is smaller than 1) you are stretching the graph vertically.

OpenStudy (anonymous):

I mean horizontally, not vertically.

OpenStudy (anonymous):

lets just do a couple of points. Tell me this. \[f(x)= \frac{1}{4x}\] when x=1/16, y=? when x=1/8, y=? when x=1/4, y=? when x=1/2, y=? when x=1, y=?

OpenStudy (anonymous):

or at leats tell me the order of the above I posted, from the greater to a smaller output.

OpenStudy (anonymous):

ohhhh wait i think i get it now thank you

OpenStudy (anonymous):

yes, so your graph would roughly look like,